OpenStudy (anonymous):

The functions f and g are defined as followed. f(x)=3x+4 g(x)=-3x^3-2 Find f(-5) and G(-3)

OpenStudy (anonymous):

for f(-5) Substitute -5 in place of x in the equation of f(x) and for g(-3) substitute -3 in place of x in g(x)

OpenStudy (anonymous):

f(-5) = -15 + 4 = -11 g(-3) = -3 (- 3)^3 -2 = -81 -2 = -83

OpenStudy (anonymous):

@ganibl when I worked out the second one with g I got 79, did I do my exponent wrong?

OpenStudy (anonymous):

No U got it right :) I did wrong :)

OpenStudy (anonymous):

-3 (-27) - 2 = 81 - 2 = 79

OpenStudy (anonymous):

Okay will you check this one for me? I'm pretty sure I got the second one rights I just want to make sure on the first one @ganibl f(x)=4x^2-x and g(x)=-4x+4 f(-3) G(4) For the first one I got 141 and for the second one I got -12

OpenStudy (anonymous):

f(-3) = 4(-3)^2 - (-3) = 4*9 + 3 = 36 + 3 = 39 g(4) = -4(4) + 4 = -16 + 4 = -12

OpenStudy (anonymous):

@ganibl oh okay I see what I did, thank you so much :)